# When Can LLMs Actually Correct Their Own Mistakes? A Critical Survey of Self-Correction of LLMs

### v1

@@ Abstract
Self-correction is an approach to improving responses from large language models (LLMs) by refining the responses using LLMs during inference. Prior work has proposed various self-correction frameworks using different sources of feedback, including self-evaluation and external feedback. However, there is still no consensus on the question of when LLMs can correct their own mistakes, as recent studies also report negative results. In this work, we critically survey broad papers and discuss the conditions required for successful self-correction. We first find that prior studies often do not define their research questions in detail and involve impractical frameworks or unfair evaluations that over-evaluate self-correction. To tackle these issues, we categorize research questions in self-correction research and provide a checklist for designing appropriate experiments. Our critical survey based on the newly categorized research questions shows that (1) no prior work demonstrates successful self-correction with feedback from prompted LLMs in general tasks, (2) self-correction works well in tasks that can use reliable external feedback, and (3) large-scale fine-tuning enables self-correction.
@@ 1 Introduction
Self-correction has emerged as a popular approach to improve responses from large language models (LLMs) by refining them using LLMs during inference (Bai et al., 2022; Madaan et al., 2023), under the hypothesis that recognizing errors is easier than avoiding them (Saunders et al., 2022). Extensive studies on self-correction have been conducted in various tasks, including arithmetic reasoning, code generation, and question answering (Gao et al., 2023; Shinn et al., 2023). The simplest approach of self-correction prompts LLMs to provide feedback on their own responses and refine the responses using the feedback (Huang et al., 2024). As in Figure 1, self-correction has also been studied using additional information for improving feedback, including external tools such as code interpreters (Chen et al., 2024c; Gou et al., 2024), external knowledge retrieved via web search (Gao et al., 2023; Jiang et al., 2023b), or fine-tuning (Welleck et al., 2023; Ye et al., 2023). However, recent studies also report negative results indicating that LLMs cannot self-correct (Huang et al., 2024; Gou et al., 2024) or even self-detect (Chen and Shu, 2024; Tyen et al., 2024; Hong et al., 2024; Kamoi et al., 2024) their own mistakes at least in certain conditions. These conflicting observations indicate that further analysis of self-correction is needed.
In this work, we provide a critical survey to investigate the conditions required for successful self-correction. First, our analysis finds that prior studies often do not define their research questions in detail. As a result, many papers fail to provide appropriate experiments to evaluate the research questions they implicitly target. To address this issue, we establish categories of research questions in self-correction research (§3.1) and discuss frameworks that should be used for verifying each research question (§3.2). Finally, we provide a checklist for designing appropriate experiments (§8).
Next, we analyze prior work to identify when LLMs can self-correct their mistakes, using the new definitions of the research questions. Our analysis highlights that the bottleneck is in the feedback generation (§7). Specifically, (1) no prior work shows successful self-correction with feedback from prompted LLMs in general tasks (§4), (2) self-correction works well in tasks where reliable external feedback is available (§5), (3) large-scale fine-tuning enables self-correction (§6), and (4) some tasks have properties exceptionally suitable for self-correction (§4). In summary, our analysis identifies the properties required for successful self-correction as follows:
@@
[RQ2] When can LLMs self-correct the best-possible initial responses with external feedback?
@@
Self-correction is often not compared with sufficiently strong baselines, and it is still unclear whether it is better than other approaches. (§3.3)
This survey is organized as follows. Section 2 provides an overview of self-correction. Section 3 introduces a new approach to classify research questions and frameworks in self-correction research. Sections 4, 5, and 6 analyze prior work in self-correction with in-context learning, external information, and fine-tuning, respectively. Section 7 summarizes our findings from the analysis. Section 8 provides a checklist for self-correction research. Section 9 explains differences from other surveys. Section 10 provides studies related to self-correction. Section 11 provides future directions.
@@
| Self-Debug (2024c) | Fair-Asym. | GPT-3.5-T, GPT-4 | – | Code Interpreter | – | Text-to-Code | – | – | – | – |
@@
| Huang et al. (2024) | Intrinsic | GPT-4-T, GPT-3.5-T | – | – | – | GSM8k | CSQA, HotpotQA | – | – | – |
@@
Self-Debug (2024c), CRITIC (2024),
@@ 2 Self-Correction of LLMs
The term “self-correction” is used in a wide range of scenarios, from a strict definition in which LLMs refine their own responses by themselves (Madaan et al., 2023; Huang et al., 2024) to broader concepts that also involve feedback from external tools or knowledge (Shinn et al., 2023; Gou et al., 2024). In this work, we define self-correction as a framework that refines responses from LLMs by using LLMs during inference, possibly using external tools or knowledge. As shown in Table 1, Figure 2, and Figure 3, self-correction has been studied in various frameworks with different sources of feedback.
@@ 2.1 Frameworks
Self-correction often consists of three stages including feedback generation (Kim et al., 2023; Madaan et al., 2023; Shinn et al., 2023; Huang et al., 2024):
@@
Cross-model correction generates feedback or refines the responses using models different from the model that generates initial responses. Cross-model correction has been mostly studied in the settings of correcting mistakes of large proprietary LLMs using small fine-tuned models (Welleck et al., 2023; Akyurek et al., 2023; Paul et al., 2024) or multi-agent debate of multiple models with similar capabilities (Liang et al., 2023; Li et al., 2023; Cohen et al., 2023b; Du et al., 2023; Zhang et al., 2023a; Chen et al., 2024b; Chan et al., 2024; Wang et al., 2024).
@@ 2.2 Sources of Feedback
External Information (§5).
Self-correction often relies on external information, including external tools such as code executors (Jiang et al., 2023a; Gou et al., 2024; Chen et al., 2024c; Stengel-Eskin et al., 2024), symbolic reasoners (Pan et al., 2023), proof assistant (First et al., 2023), or task-specific metrics (Xu et al., 2023), external knowledge from search engines (Jiang et al., 2023b; Gao et al., 2023; Zhao et al., 2023), Wikipedia (Yu et al., 2023; Zhao et al., 2023), or other corpora (Peng et al., 2023; Zhao et al., 2023), oracle information such as ground-truth answers (Kim et al., 2023; Shinn et al., 2023), human feedback (Chen et al., 2024a), or stronger models (Zhang et al., 2024).
Fine-tuning (§6).
@@ 2.3 Tasks
Self-correction has been studied in Reasoning: arithmetic reasoning (Madaan et al., 2023; Nathani et al., 2023; Gou et al., 2024), code generation (Jiang et al., 2023a; Charalambous et al., 2023; Gou et al., 2024; Chen et al., 2024c; Olausson et al., 2024), proof generation (First et al., 2023), logical reasoning (Pan et al., 2023), Knowledge: closed-book QA (Shinn et al., 2023; Gao et al., 2023; Jiang et al., 2023b; Gou et al., 2024), Context-based Generation: dialogue generation (Madaan et al., 2023; Peng et al., 2023), text summarization (Saunders et al., 2022), Open-ended Generation: conditional text generation (Ye et al., 2023; Schick et al., 2023), story generation (Yang et al., 2022b), detoxification (Schick et al., 2021; Bai et al., 2022; Gou et al., 2024; Phute et al., 2024), Others: machine translation (Chen et al., 2023b; Raunak et al., 2023; Ki and Carpuat, 2024), information retrieval (Gero et al., 2023), visual instruction following (Lee et al., 2023), and prompt optimization (Pryzant et al., 2023; Mehrabi et al., 2023; Yang et al., 2024).
@@ 2.4 Differences from Related Approaches
In this work, we define self-consistency (Wang et al., 2023) or generate-and-rank (Shen et al., 2021; Weng et al., 2023) to be different from self-correction because these approaches do not refine responses and assume that LLMs generate correct answers with a reasonable probability. We discuss these methods in Section 3.3 as strong baselines.
@@ 3 Research Questions
We provide a new approach to classify research questions and frameworks in self-correction.
@@ 3.1 RQs in Self-Correction Research
RQ Group A.
RQ Group B.
The performance of the final outputs of self-correction.
@@
Requirements.
Experiments for verifying these research questions need to satisfy different requirements, as in Table 3. Evaluation: RQ1 and RQ2 only require to show that self-correction improves performance from the initial responses. RQ3 requires comparison with strong baselines (§3.3). Initial Responses: RQ1 should be evaluated on frameworks that refine responses using the same model without additional information. RQ2 needs to be evaluated on frameworks that use the best-possible initial responses. RQ3 only focuses on the final performance, so it is not necessary to use strong initial responses.
@@ 3.2 Frameworks for Verifying RQs
Prior work often categorizes self-correction frameworks based on approaches for generating feedback (§2). However, we point out that they also need to be categorized by the quality of initial responses because the frameworks we need to use for verifying different research questions vary by whether they use the best-possible initial responses (§3.1).
We propose categories of (same-model) self-correction that correspond to different research questions (§3.1), as shown in Figure 2.
@@
In this work, we focus on categorizing self-correction frameworks that do not involve multiple language models with different architectures. Cross-model correction uses different models for initial response generation and self-correction, so it is unsuitable for evaluating whether LLMs can improve their own initial responses [RQ1, RQ2]. However, we note that it can be used to evaluate [RQ3] whether the final responses from self-correction are better than other methods.
@@
Realistic frameworks can be categorized by whether they use the best-possible initial responses. Fair self-correction represents frameworks that refine the best-possible initial responses. (1) Intrinsic self-correction (Huang et al., 2024) uses the same model and information for initial response generation and self-correction. This framework can assess [RQ1] whether LLMs can self-correct based solely on their inherent capabilities. (2) Fair-asymmetric self-correction uses additional information for self-correction, but also uses information to improve initial response generation as much as possible. For example, self-correction with code interpreters (Chen et al., 2024c; Gou et al., 2024) is not intrinsic but fair because we cannot easily use code interpreters to directly improve the initial response generation. This framework can evaluate [RQ2] whether LLMs can self-correct the best-possible initial responses using external feedback. Unfair self-correction (or unfair-asymmetric self-correction) represents frameworks that are practical but do not use the best-possible initial responses. For example, methods that use search engines only for self-correction (Gao et al., 2023; Yu et al., 2023) are unfair because they can use search engines to directly improve the initial response generation. Unfair self-correction can evaluate [RQ3] whether the final responses from self-correction outperform other methods but cannot evaluate [RQ2] whether self-correction can improve the best-possible initial responses.
@@
3.3 Strong Baselines
Self-correction involves multiple LLM calls for generating feedback and refinement. Therefore, to claim that [RQ3] the performance of the final outputs from self-correction frameworks is better than other approaches, it should be compared with sufficiently strong baselines, possibly relying on additional computation. Many studies do not compare their methods with strong baselines, although some studies pointed out this issue by comparing self-correction with self-consistency (Gou et al., 2024; Huang et al., 2024) or pass@k in code generation (Zhang et al., 2023b; Olausson et al., 2024).
Self-Consistency
(Wang et al., 2023) generates multiple responses for the same input and takes the majority vote of the final answers in reasoning tasks. The idea of selecting good responses using consistency between multiple responses has been extended to other tasks such as text generation (Manakul et al., 2023; Elaraby et al., 2023; Chen et al., 2023c) and code generation (Shi et al., 2022).
Generate-and-Rank
is an approach to generating multiple responses and selecting the best response using verifiers. Post-hoc approach ranks responses using self-evaluation (Weng et al., 2023; Zhang et al., 2023d), confidence (Manakul et al., 2023), fine-tuned verifiers (Cobbe et al., 2021; Shen et al., 2021; Lightman et al., 2024), or verifiers with external tools (Shi et al., 2022; Chen et al., 2023a; Ni et al., 2023). Feedback-guided decoding generates multiple responses and selects the best response for each reasoning step using generation probability (Hao et al., 2023; Tyen et al., 2024), prompted self-evaluation (Jung et al., 2022; Creswell and Shanahan, 2022; Xie et al., 2023; Yao et al., 2023; Miao et al., 2024), or fine-tuned verifiers (Uesato et al., 2022; Tafjord et al., 2022; Yang et al., 2022a; Asai et al., 2024).
@@ 4 Self-Correction with Prompting
However, recent studies report that intrinsic self-correction does not improve or even degrade the performance in tasks such as arithmetic reasoning, closed-book QA (Huang et al., 2024; Gou et al., 2024), code generation (Gou et al., 2024; Olausson et al., 2024), plan generation (Valmeekam et al., 2023), and graph coloring (Stechly et al., 2023). Several studies claim that a bottleneck is in the feedback generation, and it is difficult to generate reliable feedback on their responses only by prompting themselves (Gou et al., 2024; Huang et al., 2024; Olausson et al., 2024).
Unfair Settings.
The conflicting positive and negative results motivate us to analyze when LLMs can self-correct only by prompting themselves. Specifically, we assess whether prior studies satisfy the requirements to verify that [RQ1] LLMs can self-correct their responses based solely on their inherent capabilities. As in Table 4, we find that many studies use either oracle information in the self-correction processes (unrealistic frameworks) or weak prompts for generating initial responses (unfair settings), which over-evaluate self-correction. Consequently, we conclude that no major work shows successful self-correction of responses from LLMs using feedback generated by prompting themselves under fair settings in general tasks. Oracle Information: RCI Prompting (Kim et al., 2023) uses ground-truth answers and does not apply self-correction when the initial responses are correct, which unfairly ignores mistakes of incorrectly updating correct responses. Reflexion (Shinn et al., 2023) generates feedback by using an exact match between the generated and ground-truth answers, which cannot be accessed in real-world applications. Weak Initial Responses: Bai et al. (2022) propose self-correction for detoxifying harmful responses, but initial response generation is not instructed to generate harmless responses. Although detecting harmful contents using LLMs is a reasonable approach, this setting is not the self-correction from best-possible initial responses. Self-Refine (Madaan et al., 2023) uses instructions or few-shot examples that do not correctly correspond to the target task only for initial response generation, while using appropriate instructions for self-correction, as shown in Table 9 and 10. These settings evaluate improvement from weak initial responses, which over-evaluate self-correction.
@@
Although prior studies show that intrinsic self-correction is generally difficult, some tasks have properties that enable self-correction. First, tasks with decomposable responses, whose feedback generation can be reduced into easier tasks, allow to generate reasonable feedback only by prompting LLMs. For example, CoVe (Dhuliawala et al., 2023) targets tasks of generating multiple answers, such as Who are some politicians who were born in Boston? The feedback generation can be decomposed into easier sub-tasks of verifying each answer. Second, verifiable tasks, in which the correctness of the responses can be verified easily without external information, are suitable for self-correction. For example, tree-of-thought (Yao et al., 2023) is a generate-and-rank method for verifiable tasks, such as Game of 24, a task to find arithmetic operations to obtain 24 using provided four integers. We expect self-correction to work well on these tasks because the situation is identical to using the oracle information to generate feedback (e.g., the target answer is 24). These tasks are exceptionally suitable for self-correction, but many real-world tasks do not satisfy these properties.
@@
| Self-Debug (2024c) | Text-to-Code | – | Code Interpreter |
@@
5 Self-Correction with External Tools or Knowledge
Given the observation that feedback generation is a bottleneck of self-correction (§4), improving feedback using external tools or knowledge is a promising direction. External tools used for self-correction include code interpreters for code generation tasks (Chen et al., 2024c; Gou et al., 2024) and symbolic reasoners for logical reasoning tasks (Pan et al., 2023). A popular source of knowledge is search engines, which are often used with queries generated from initial responses to retrieve information for validating their correctness (Gao et al., 2023; Jiang et al., 2023b). Prior work widely agrees that self-correction can improve LLM responses when reliable external tools or knowledge suitable for improving feedback are available [RQ2].

### v2

@@ Abstract
Self-correction is an approach to improving responses from large language models (LLMs) by refining the responses using LLMs during inference. Prior work has proposed various self-correction frameworks using different sources of feedback, including self-evaluation and external feedback. However, there is still no consensus on the question of when LLMs can correct their own mistakes, as recent studies also report negative results. In this work, we critically survey broad papers and discuss the conditions required for successful self-correction. We first find that prior studies often do not define their research questions in detail and involve impractical frameworks or unfair evaluations that over-evaluate self-correction. To tackle these issues, we categorize research questions in self-correction research and provide a checklist for designing appropriate experiments. Our critical survey based on the newly categorized research questions shows that (1) no prior work demonstrates successful self-correction with feedback from prompted LLMs, except for studies in tasks that are exceptionally suited for self-correction, (2) self-correction works well in tasks that can use reliable external feedback, and (3) large-scale fine-tuning enables self-correction.
@@ 1 Introduction
Self-correction is a popular approach to improve responses from large language models (LLMs) by refining them using LLMs during inference (Bai et al., 2022; Madaan et al., 2023). Extensive studies on self-correction have been conducted in various tasks, including arithmetic reasoning, code generation, and question answering (Gao et al., 2023; Shinn et al., 2023). The simplest approach of self-correction prompts LLMs to provide feedback on their own responses and refine the responses using the feedback (Huang et al., 2024a), under the hypothesis that recognizing errors is easier than avoiding them (Saunders et al., 2022). As in Figure 1, self-correction has also been studied using additional information for improving feedback, including external tools such as code interpreters (Chen et al., 2024d; Gou et al., 2024), external knowledge retrieved via web search (Gao et al., 2023; Jiang et al., 2023b), or fine-tuning (Welleck et al., 2023; Ye et al., 2023). However, recent studies also report negative results indicating that LLMs cannot self-correct (Huang et al., 2024a; Gou et al., 2024; Li et al., 2024b) or even self-detect (Chen and Shu, 2024; Tyen et al., 2024; Hong et al., 2024; Jiang et al., 2024; Kamoi et al., 2024) their own mistakes at least in certain conditions. These conflicting observations indicate that further analysis of self-correction is needed.
In this work, we provide a critical survey to investigate the conditions required for successful self-correction. First, our analysis finds that prior studies often do not define their research questions in detail. As a result, many papers fail to provide appropriate experiments to evaluate the research questions they implicitly target. To address this issue, we categorize research questions in self-correction research (§3.1) and discuss frameworks that should be used for verifying each research question (§3.2). Finally, we provide a checklist for designing appropriate experiments (§8).
Next, we analyze prior work to identify when LLMs can self-correct their mistakes, using the new definitions of the research questions. Our analysis highlights that the bottleneck is in the feedback generation (§7). Specifically, (1) no prior work shows successful self-correction with feedback from prompted LLMs in general tasks (§4), (2) self-correction works well in tasks where reliable external feedback is available (§5.1), (3) large-scale fine-tuning enables self-correction (§5.2), and (4) some tasks have properties exceptionally suitable for self-correction (§4). In summary, our analysis identifies the properties required for successful self-correction as follows:
@@
In general tasks, no prior work shows reliable evidence of successful self-correction with in-context learning. (§4)
In tasks with specific properties that are exceptionally favorable for self-correction (e.g., responses are decomposable), self-correction is effective even with in-context learning. (§4)
[RQ2] When can LLMs self-correct the best-possible initial responses with external information?
@@
Self-correction is often not compared with sufficiently strong baselines, and it is still unclear whether it is better than other approaches. (§6)
This survey is organized as follows. Section 2 provides an overview of self-correction. Section 3 introduces a new approach to classify research questions and frameworks in self-correction research. Sections 4 and 5 analyze prior work in self-correction with in-context learning and external information (external tools, external knowledge, fine-tuning), respectively. Section 6 explains related approaches that should be compared with self-correction as baselines. Section 7 summarizes our findings from the analysis. Section 8 provides a checklist for self-correction research. Section 9 explains differences from other surveys. Section 10 provides studies related to self-correction. Section 11 provides future directions.
@@
| Self-Debug (2024d) | Fair-Asym. | GPT-3.5-T, GPT-4 | – | Code Interpreter | – | Text-to-Code | – | – | – | – |
@@
| Huang et al. (2024a) | Intrinsic | GPT-4-T, GPT-3.5-T | – | – | – | GSM8k | CSQA, HotpotQA | – | – | – |
@@
Self-Debug (2024d), CRITIC (2024),
@@ 2 Self-Correction of LLMs
The term “self-correction” is used in a wide range of scenarios, from a strict definition in which LLMs refine their own responses by themselves (Madaan et al., 2023; Huang et al., 2024a) to broader concepts that also involve feedback from external tools or knowledge (Shinn et al., 2023; Gou et al., 2024). In this work, we define self-correction as a framework that refines responses from LLMs using LLMs during inference, possibly with external tools or knowledge. As in Table 1, Figure 2, and Figure 3, self-correction has been studied in various frameworks with different sources of feedback.
@@ 2.1 Frameworks
Prior studies propose self-correction frameworks with various different architectures.
@@
Self-correction often consists of three stages including feedback generation (Kim et al., 2023; Madaan et al., 2023; Shinn et al., 2023; Huang et al., 2024a):
@@
Cross-model correction generates feedback or refines the responses using models different from the model that generates initial responses. Cross-model correction has been mostly studied in the settings of correcting mistakes of large proprietary LLMs using small fine-tuned models (Welleck et al., 2023; Akyurek et al., 2023; Paul et al., 2024) or multi-agent debate of multiple models with similar capabilities (Liang et al., 2023; Li et al., 2023; Cohen et al., 2023; Du et al., 2023; Zhang et al., 2023a; Chen et al., 2024b; Chan et al., 2024; Wang et al., 2024a).
@@ 2.2 Sources of Feedback
External Information (§5.1).
Self-correction often relies on external information, including external tools such as code executors (Jiang et al., 2023a; Gou et al., 2024; Chen et al., 2024d; Stengel-Eskin et al., 2024), symbolic reasoners (Pan et al., 2023), proof assistant (First et al., 2023), or task-specific metrics (Xu et al., 2023), external knowledge from search engines (Jiang et al., 2023b; Gao et al., 2023; Zhao et al., 2023), Wikipedia (Yu et al., 2023; Zhao et al., 2023), or other corpora (Peng et al., 2023; Zhao et al., 2023), oracle information such as ground-truth answers (Kim et al., 2023; Shinn et al., 2023), human feedback (Chen et al., 2024a), or stronger models (Zhang et al., 2024).
Fine-tuning (§5.2).
@@ 2.3 Tasks
Self-correction has been studied in various tasks, including Reasoning: arithmetic reasoning (Madaan et al., 2023; Nathani et al., 2023; Gou et al., 2024), code generation (Jiang et al., 2023a; Charalambous et al., 2023; Gou et al., 2024; Chen et al., 2024d; Olausson et al., 2024), proof generation (First et al., 2023), logical reasoning (Pan et al., 2023), Knowledge: closed-book QA (Shinn et al., 2023; Gao et al., 2023; Jiang et al., 2023b; Gou et al., 2024), Context-based Generation: dialogue generation (Madaan et al., 2023; Peng et al., 2023), text summarization (Saunders et al., 2022), Open-ended Generation: conditional text generation (Ye et al., 2023; Schick et al., 2023), story generation (Yang et al., 2022b), detoxification (Schick et al., 2021; Bai et al., 2022; Gou et al., 2024; Phute et al., 2024), Others: machine translation (Chen et al., 2023b; Raunak et al., 2023; Ki and Carpuat, 2024), information retrieval (Gero et al., 2023), vision language tasks (Yin et al., 2023; Ge et al., 2023; Zhou et al., 2024; Lee et al., 2024; Huang et al., 2024b; Liu et al., 2024), and prompt optimization (Pryzant et al., 2023; Mehrabi et al., 2023; Yang et al., 2024).
@@ 2.4 Differences from Related Approaches
In this work, we define self-consistency (Wang et al., 2023) or generate-and-rank (Shen et al., 2021; Weng et al., 2023) to be different from self-correction because these approaches do not refine responses and assume that LLMs generate correct answers with a reasonable probability. We discuss these methods in Section 6 as strong baselines that should be compared with self-correction.
@@ 3 Research Questions
We find that prior studies often do not define their research questions in detail and fail to use appropriate self-correction frameworks in their experiments. We propose a new approach to classify research questions and frameworks in self-correction.
@@ 3.1 RQs in Self-Correction Research
Requirements for Verifying RQs.
Experiments for verifying these research questions need to satisfy different requirements, as shown in Table 3. External Information: RQ1 needs to be evaluated on frameworks that refine responses using the same model without additional information. RQ2 and RQ3 can be evaluated on frameworks that use external information. Initial Responses: RQ1 and RQ2 need to be evaluated on frameworks that use the best-possible initial responses. RQ3 is about the final performance, so it is not necessary to start from strong initial responses. Evaluation: RQ1 and RQ2 only require to show that self-correction improves performance from the initial responses. RQ3 requires comparison with strong baselines (§6).
@@ 3.2 Frameworks for Verifying RQs
Prior work often categorizes self-correction frameworks based on approaches for generating feedback (§2). However, we point out that we also need to categorize them by the quality of initial responses because the frameworks we need to use for verifying different research questions vary by whether they use the best-possible initial responses (§3.1).
We propose categories of (same-model) self-correction that correspond to different research questions (§3.1), as shown in Figure 2. Specifically, we propose to categorize the self-correction frameworks as follows.
@@
In this work, we focus on categorizing self-correction frameworks that do not involve multiple language models with different architectures. Cross-model correction uses different models for initial response generation and self-correction, so it is unsuitable for evaluating whether LLMs can improve their own initial responses [RQ1, RQ2]. However, it can be used to evaluate [RQ3] whether the final responses from self-correction are better than other methods.
@@
Realistic frameworks can be categorized by whether they use the best-possible initial responses. Fair self-correction represents frameworks that refine the best-possible initial responses. (1) Intrinsic self-correction (Huang et al., 2024a) uses the same model and information for initial response generation and self-correction. Intrinsic self-correction can be used to assess [RQ1] whether LLMs can self-correct based solely on their inherent capabilities. (2) Fair-asymmetric self-correction uses additional information for self-correction, but also uses information to improve initial response generation as much as possible. For example, self-correction with code interpreters (Chen et al., 2024d; Gou et al., 2024) is not intrinsic but fair because we cannot easily use code interpreters to directly improve the initial response generation. Fair-asymmetric self-correction can be used to evaluate [RQ2] whether LLMs can self-correct the best-possible initial responses using external information. Unfair self-correction (or unfair-asymmetric self-correction) represents frameworks that are practical but do not use the best-possible initial responses. For example, methods that use search engines only for self-correction (Gao et al., 2023; Yu et al., 2023) are unfair because they can use search engines to directly improve the initial response generation. Unfair self-correction can evaluate [RQ3] whether the final responses from self-correction outperform other methods but cannot evaluate [RQ2] whether self-correction can improve the best-possible initial responses.
@@ 4 Self-Correction with Prompting
[RQ1] Can LLMs self-correct their best-possible initial responses based solely on the inherent capabilities?
@@
However, recent studies report that intrinsic self-correction does not improve or even degrade the performance in tasks such as arithmetic reasoning, closed-book QA (Huang et al., 2024a; Gou et al., 2024), code generation (Gou et al., 2024; Olausson et al., 2024), plan generation (Valmeekam et al., 2023), and graph coloring (Stechly et al., 2023). Several studies claim that a bottleneck is in the feedback generation, and it is difficult to generate reliable feedback on their responses only by prompting themselves (Gou et al., 2024; Huang et al., 2024a; Olausson et al., 2024).
Unrealistic or Unfair Settings.
The conflicting positive and negative results motivate us to analyze when LLMs can self-correct only by prompting themselves. Specifically, we assess whether prior studies satisfy the requirements to verify that [RQ1] LLMs can self-correct their responses based solely on their inherent capabilities. As in Table 4, we find that many studies use either oracle information in the self-correction processes (unrealistic frameworks) or weak prompts that can be easily improved for generating initial responses (unfair settings), which over-evaluate self-correction. Consequently, we conclude that no major work shows successful self-correction of responses from LLMs using feedback generated by prompting themselves under fair settings in general tasks. Oracle Information: RCI Prompting (Kim et al., 2023) uses ground-truth answers and does not apply self-correction when the initial responses are correct, which unfairly ignores mistakes caused by updating correct responses incorrectly. Reflexion (Shinn et al., 2023) generates feedback by using an exact match between the generated and ground-truth answers, which cannot be accessed in real-world applications. Weak Initial Responses: Detoxifying harmful responses is a popular task in self-correction research, but prior studies often study in situations where initial response generation is not instructed to generate harmless responses (Bai et al., 2022; Wang et al., 2024b). Although detecting harmful contents using LLMs is a reasonable research topic, this setting is not the self-correction from best-possible initial responses, since we can improve the initial response generation process by instructing not to generate harmful responses. As more obvious weak prompts, Self-Refine (Madaan et al., 2023) uses instructions or few-shot examples that do not correctly correspond to the target task only for initial response generation (e.g., providing wrong target labels in few-shot examples), while using appropriate instructions for self-correction, as shown in Table 9 and 10. These settings evaluate improvement from weak initial responses, which over-evaluate the improvement by self-correction.
@@
Although our analysis of prior studies shows that intrinsic self-correction is difficult in general, some tasks have properties that make feedback generation easy and enable intrinsic self-correction. For example, CoVe (Dhuliawala et al., 2023) is an intrinsic self-correction method for tasks of generating multiple answers, such as Name some politicians who were born in NY, New York. Generated responses include multiple answers, but the feedback generation can be decomposed into easier sub-tasks of verifying each answer. Tasks with decomposable responses are one of the few groups of tasks for which verification is clearly easier than generation, which enables intrinsic self-correction. However, many real-world tasks do not satisfy this property.
@@
| Self-Debug (2024d) | Text-to-Code | – | Code Interpreter |
@@
5 Self-Correction with External Information
[RQ2] Can LLMs self-correct their best-possible initial responses assisted by external information?
This section analyzes self-correction frameworks that make use of external tools, external knowledge, and fine-tuning.
5.1 Self-Correction with External Tools or Knowledge
Given the observation that feedback generation is a bottleneck of self-correction (§4), improving feedback using external tools or knowledge is a promising direction. External tools used for self-correction include code interpreters for code generation tasks (Chen et al., 2024d; Gou et al., 2024) and symbolic reasoners for logical reasoning tasks (Pan et al., 2023). A popular source of knowledge is search engines, which are often used with queries generated from initial responses to retrieve information for validating their correctness (Gao et al., 2023; Jiang et al., 2023b). These prior studies widely agree that self-correction can improve LLM responses when reliable external tools or knowledge suitable for improving feedback are available.